(1+x+x^2)^171 what is the sum of the coefficients?

Grade:12

3 Answers

Ravi
9 years ago
Put x=1 and solve the expression to get the sum of coefficients.
(1+1+1)^171
.................................................................................................................................................................................................................................................................................
Shobhit Varshney IIT Roorkee
9 years ago
Hi,

If we expand, (1+x+x2)171= ao+ a1x+ a2x2….... (it would be a function of x.)
So, To get the sum of the coefficients,
put x = 1 in the above equation,
(1+1+1)171= ao+ a1+ a2+ ….................... = 3171

thanks
rahul
9 years ago
the answer will be 3^171